# The dimension and Bose distance of certain primitive BCH codes

### v2

@@ Abstract
Bose-Ray-Chaudhuri-Hocquenghem (BCH) codes are a significant class of cyclic codes that play an important role in both theoretical research and practical applications. Their strong error-correcting abilities and efficient encoding and decoding methods make BCH codes widely applicable in various areas, including communication systems, data storage devices, and consumer electronics. Although BCH codes have been extensively studied, the parameters of BCH codes are not known in general. Let $q$ be a prime power and $m$ be a positive integer. Denote by $\mathcal{C}_{(q,m,\delta)}$ the narrow-sense primitive BCH code with length $q^{m}-1$ and designed distance $\delta$. As of now, the dimensions of $\mathcal{C}_{(q,m,\delta)}$ are fully understood only for $m\leq 2$. For $m\geq 4$, the dimensions of $\mathcal{C}_{(q,m,\delta)}$ are known only for the range $2\leq\delta\leq q^{\lfloor(m+1)/2\rfloor+1}$ and for a limited number of special cases. In this paper, we determined the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}].$ Additionally, we have also extended our results to primitive BCH codes that are not necessarily narrow-sense.
@@ I Introduction
BCH codes have been extensively studied in the literature [1, 2, 3, 8, 10, 16, 13, 20, 23, 29, 28, 26, 27, 30, 31, 34, 42, 48, 9, 45, 19, 44, 46, 18, 43, 4, 14, 15, 39, 49]. Despite decades of research, our understanding of the dimension and minimum distance of BCH codes remains limited. As noted by Charpin [11], it is a challenging problem to determine the minimum distance $d$ of BCH codes in general. However, the BCH bound [24, 7] indicates that $d\geq d_{B}$. The equality can even be obtained for some codes. For example, the narrow sense BCH code $C_{(q,m,\delta)}$ satisfy $d=d_{B}$ if $\delta$ divides $q^{m}-1$[35, p. 259]. Furthermore, it has been conjectured that $d\leq d_{B}+4$ for a narrow-sense primitive BCH code. Therefore, determining the Bose distance is also valuable for us to understand BCH codes. Readers may refer to [4, 21, 14, 15, 16, 12] for some previous works on the Bose distance of $\mathcal{C}_{(q,m,\delta)}$. Determining the dimension of BCH codes is also difficult in general [17, 39]. For narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$, the dimension is fully determined for $m\leq 2$ [31, 39]. For $m\geq 3$, the problem remains largely unsolved. For convenience, we summarize some known results regarding the dimension of BCH codes of $\mathcal{C}_{(q,m,\delta)}$ in Table I. For primitive BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ that are not necessarily narrow-sense, there are quite limited results available. See [33] for the dimensions of some BCH codes that are not narrow-sense. The reader may also refer to [40, Table 1] and [17] for summaries of the existing results on the parameters of BCH codes.
@@
It is well known that the dimension and Bose distance of primitive BCH codes are closely related to $q$-cyclotomic coset modulo $q^{m}-1$. Specifically, the dimension of BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ can be given by calculating the sum of the sizes of distinct $q$-cyclotomic cosets $C_{a}$ modulo $q^{m}-1$ for all $a\in[b,b+\delta-2]$. Therefore, to determine the dimension of the primitive BCH code $\mathcal{C}_{(q,m,\delta,b)}$, it suffices to identify the size of $q$-cyclotomic coset $C_{a}$ for $a\in[b,b+\delta-2]$, and enumurate distinct cosets correspond to each size. Note that the size of a coset must be a divisor of $m$, which makes determining the size of any given coset relatively straightforward. However, finding the number of distinct cosets requires more effort. Yue et al. [48] introduced the concept of the coset leader, defined as the smallest integer in each coset. As a result, the number of distinct cosets corresponds exactly to the number of coset leaders. Moreover, determining the Bose distance of $C_{(q,m,\delta,b)}$ reduces to identifying the minimal coset leader greater than $\delta+b-2$.
@@
The $q$-adic expansion of the integer $a\in[0,q^{m}-1]$ provides a convenient way for determining whether $a$ is the coset leader of $C_{a}$, which is a technique widely employed in the study of BCH codes. However, determining the total number of coset leaders in $[b,b+\delta-2]$ is very difficult in general due to their highly irregular distribution. Only for sufficiently small $a$, the integer $a$ is a coset leader if and only if $q$ does not divide $a$. Consequently, the number of coset leaders equals the number of integers not divisible by $q$ in the range $[b,b+\delta-2]$ when $b+\delta-2$ is small. This observation was applied in [48] to determine the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $12\leq\delta\leq q^{\lfloor(m+1)/2\rfloor}+$ 1, and for $q^{m/2}+2\leq\delta\leq 2q^{m/2}+1$ when $m$ is even. Later, Liu et al. [33] improved this result by giving the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}.$ Beyond this range, the dimension is determined for a few special cases. For example, Mann [36] established the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=q^{t}$. Cherchem et al. [12] give the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=\frac{a(q^{m}-1)}{q-1}$ and $\delta=aq^{m}-1$ with $a\in[1,q-1]$. By identifying the first few largest coset leaders and the sizes of corresponding cosets, the parameters of $C_{(q,m,\delta)}$ for $\delta$ as the four largest coset leaders were determined in [16] [32] and [38]. One can also obtain the parameters of some BCH codes by comparing them with those of codes with known parameters. Building on Mann’s work [36], Ding et al. determined the dimensions of $\mathcal{C}_{(q,m,q^{t}-1)}$ [14, Theorem 14], $\mathcal{C}_{(q,m,q^{t}+b)}$ for $11\leq b\leq\lfloor(q^{t}-1)/q^{m-t}\rfloor+$ 1 [15, Theorem 13], and $\mathcal{C}_{(q,m,q^{t}-2)}$ for $q>2$ [14, Theorem 15]. Recently, the dimension of $\mathcal{C}_{(q,m,q^{t}+b)}$ for $\lceil{m}/{2}\rceil\leq t<m$ and $0\leq b<q^{m-t}+\sum\limits_{i=1}^{\lfloor{t}/{s}\rfloor-1}q^{m+i(m-t)}$, and $\mathcal{C}_{(q,m,q^{t}-b)}$ for $\lceil{m}/{2}\rceil<t<m$ and $0\leq b<(q-1)\sum\limits_{i=1}^{m-t}q^{i}$ were provided in [21].
@@
In this paper, we determine the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/{3}\rfloor+1}.$ We also extend these results to BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ for $m\geq 4$ and positive integers $b$ and $\delta$ with $b+\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}+1$. In the existing literature, for $m\geq 4$, the dimension of $\mathcal{C}_{(q,m,\delta)}$ is known only for $\delta\in[2,q^{\lfloor(m+1)/{2}\rfloor+1}]$ and for some special cases. Moreover, it is clear that $q^{\lfloor(2m-1)/{3}\rfloor+1}\geq q^{\lfloor(m+1)/{2}\rfloor+1}\cdot q^{\lfloor(m-4)/{6}\rfloor}$. Therefore, our results provide the dimension of $C_{(q,m,\delta)}$ for $\delta$ in a much larger range than previously known.
@@
Our results rely critically on the analysis of the distribution of coset leaders in $[1,q^{\lfloor(2m-1)/3\rfloor+1}]$. Note that an integer $a\in[0,q^{m}-1]$ cannot be a coset leader if $q$ divides $a$. Therefore, to determine the number of coset leaders within a given range, we only need to count the integers that are neither divisible by $q$ nor coset leaders. We partitioned such integers in $\left[1,q^{\left\lfloor(2m-1)/3\right\rfloor+1}\right]$ into distinct classes with uniform representations by applying a one-to-one correspondence between the integer in $[0,q^{m}-1]$ and length-$m$ sequences of non-negative integers less than $q$. This classification allows us to compute the number of integers in each class through a uniform method. Furthermore, we divide the interval $[1,\delta-1]$ into subintervals based on the distribution of coset leaders, so that we can establish an equation to express the number of integers neither divisible by $q$ nor coset leaders in each subinterval. Consequently, we derive the exact count of coset leaders in $[1,\delta-1]$ for each integer $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$. Combining this with our findings on the sizes of cosets, we ultimately determined the dimension of the code $\mathcal{C}_{(q,m,\delta)}$ for each integer $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$. On the other hand, the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ is indeed equal to the smallest integer not less than $\delta$. Therefore, it is not difficult to determine the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ based on our analysis of the distribution of coset leaders.
@@
$m$ is odd $12\leq\delta\leq q^{({m+1})/{2}}+$ 1 1996 [48]
$m$ is even $12\leq\delta\leq 2q^{{m}/{2}}+$ 1 1996 [48]
$m\geq 1$ $\delta=(q-l_{0})q^{m-l_{1}-1}-1$ with $20\leq l_{0}\leq q-$ 2 and $10\leq l_{1}\leq m-$ 1 2015 [14]
$m$ is even $\delta=k(q^{{m}/{2}}+1)$ with $11\leq k\leq q-$ 1 2015 [14]
@@
$m\geq 1$ $\delta=q^{t}+b$ with $11\leq b\leq\lfloor(q^{t}-1)/q^{m-t}\rfloor+$ 1 2015 [15]
@@
$m=2$ $22\leq\delta\leq q^{m}-$ 2 2017 [33]
$m>1$ $\delta=a({q^{m}-1})/{q-1}$ or $\delta=aq^{m-1}-1$ with $11\leq a\leq q-$ 1 2020 [12]
@@
$m>1$ $\delta=q^{t}+b$ with $\lceil{m}/{2}\rceil\leq t<m$ and $0\leq b<q^{m-t}+\sum\limits_{i=1}^{\lfloor{t}/{s}\rfloor-1}q^{m+i(m-t)};$ $\delta=q^{t}-b$ with $\lceil{m}/{2}\rceil<t<m$ and $0\leq b<(q-1)\sum\limits_{i=1}^{m-t}q^{i};$ 2024 [21]
This paper is organized as follows. In Section II, we present some preliminaries that reveal the close relationship between $q$-cyclotomic cosets modulo $q^{m}-1$ and the dimension and Bose distance of BCH codes. Section III introduces some new results on $q$-cyclotomic cosets. These findings are then utilized to determine the dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor(2m-1)/{3}\rfloor+1}$ in Section IV. The result concerning the dimension of narrow-sense primitive BCH codes is subsequently extended to primitive BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ for any positive integers $\delta$ and $b$ with integers $\delta+b\leq q^{\lfloor(2m-1)/{3}\rfloor+1}+1$ in Section V. Moreover, we also apply the findings on $q$-cyclotomic cosets to obtain the Bose distance for primitive BCH codes in Section VI. We then provide examples of the BCH codes studied in this paper and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, in Section VII, as an illustration of our main results, we present the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=aq^{h+k}+b$ with integers $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, $a\in[1,q-1]$ and $b\in\left[1,q^{m-h-k}\right]$, where $h=\lfloor m/2\rfloor$. Finally, the conclusion of this paper is given in Section VIII.

### v3

@@ Abstract
BCH codes are a significant class of cyclic codes that play an important role in both theoretical research and practical applications. Their strong error-correcting abilities and efficient encoding and decoding methods make BCH codes widely applicable in various areas, including communication systems, data storage devices, and consumer electronics. Although BCH codes have been extensively studied, the parameters of BCH codes are not known in general. Let $q$ be a prime power and $m$ be a positive integer. Denote by $\mathcal{C}_{\left(q,m,\delta)\right)}$ the narrow-sense primitive BCH code with length $q^{m}-1$ and designed distance $\delta$. As of now, the dimensions of $\mathcal{C}_{(q,m,\delta)}$ are fully understood only for $m\leq 2$. For $m\geq 4$, the dimensions of $\mathcal{C}_{(q,m,\delta)}$ are known only for the range $2\leq\delta\leq q^{\lfloor(m+1)/2\rfloor+1}$ and for a limited number of special cases. In this paper, we determined the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}].$ Additionally, we have also extended our results to some primitive BCH codes that are not necessarily narrow-sense.
@@ I Introduction
BCH codes have been extensively studied in the literature [1, 2, 3, 8, 10, 16, 13, 20, 23, 29, 28, 26, 27, 30, 31, 34, 42, 48, 9, 45, 19, 44, 46, 18, 43, 4, 14, 15, 39, 49]. Despite decades of research, our understanding of the dimension and minimum distance of BCH codes remains limited. As noted by Charpin [11], it is a challenging problem to determine the minimum distance $d$ of BCH codes in general. However, the BCH bound [24, 7] indicates that $d\geq d_{B}$. The equality can even be obtained for some codes. For example, the narrow sense BCH code $\mathcal{C}_{(q,m,\delta)}$ satisfy $d=d_{B}$ if $\delta$ divides $q^{m}-1$[35, p. 259]. Furthermore, it has been conjectured that $d\leq d_{B}+4$ for a narrow-sense primitive BCH code. Therefore, determining the Bose distance is also valuable for us to understand BCH codes. Readers may refer to [4, 21, 14, 15, 16, 12] for some previous works on the Bose distance of $\mathcal{C}_{(q,m,\delta)}$. Determining the dimension of BCH codes is also difficult in general [17, 39]. For narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$, the dimension is fully determined for $m\leq 2$ [31, 39]. For $m\geq 3$, the problem remains largely unsolved. For convenience, we summarize some known results regarding the dimension of BCH codes of $\mathcal{C}_{(q,m,\delta)}$ in Table I. For primitive BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ that are not necessarily narrow-sense, there are quite limited results available. See [33] for the dimensions of some BCH codes that are not narrow-sense. The reader may also refer to [40, Table 1] and [17] for summaries of the existing results on the parameters of BCH codes.
@@
It is well known that the dimension and Bose distance of primitive BCH codes are closely related to $q$-cyclotomic coset modulo $q^{m}-1$. Specifically, the dimension of BCH codes $\mathcal{C}_{(q,m,\delta)}$ can be given by calculating the sum of the sizes of distinct $q$-cyclotomic cosets $C_{a}$ modulo $q^{m}-1$ for all $a\in[1,\delta-1]$. Therefore, to determine the dimension of the primitive BCH code $\mathcal{C}_{(q,m,\delta)}$, it suffices to identify the size of $q$-cyclotomic coset $C_{a}$ for $a\in[1,\delta-1]$, and enumurate distinct cosets correspond to each size. Note that the size of a coset must be a divisor of $m$, which makes determining the size of any given coset relatively straightforward. However, finding the number of distinct cosets requires more effort. Yue et al. [48] introduced the concept of the coset leader, defined as the smallest integer in each coset. As a result, the number of distinct cosets corresponds exactly to the number of coset leaders. Moreover, determining the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ reduces to identifying the minimal coset leader greater than $\delta-1$.
@@
The $q$-adic expansion of the integer $a\in[0,q^{m}-1]$ provides a convenient way for determining whether $a$ is the coset leader of $C_{a}$, which is a technique widely employed in the study of BCH codes. However, determining the total number of coset leaders in $b,b+\delta-2]$ is very difficult in general due to their highly irregular distribution. Only for sufficiently small $a$, the integer $a$ is a coset leader if and only if $q$ does not divide $a$. Consequently, the number of coset leaders equals the number of integers not divisible by $q$ in the range $b,b+\delta-2]$ when $b+\delta-2$ is small. This observation was applied in [48] to determine the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor(m+1)/2\rfloor}+1$, and for $q^{m/2}+2\leq\delta\leq 2q^{m/2}+1$ when $m$ is even. Later, Liu et al. [33] improved this result by giving the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}.$ Beyond this range, the dimension is determined for a few special cases. For example, Mann [36] established the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=q^{t}$. Cherchem et al. [12] give the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=\frac{a(q^{m}-1)}{q-1}$ and $\delta=aq^{m}-1$ with $a\in[1,q-1]$. By identifying the first few largest coset leaders and the sizes of corresponding cosets, the parameters of $C_{(q,m,\delta)}$ for $\delta$ as the four largest coset leaders were determined in [16] [32] and [38]. One can also obtain the parameters of some BCH codes by comparing them with those of codes with known parameters. Building on Mann’s work [36], Ding et al. determined the dimensions of $\mathcal{C}_{(q,m,q^{t}-1)}$ [14, Theorem 14], $\mathcal{C}_{(q,m,q^{t}+b)}$ for $1\leq b\leq\lfloor(q^{t}-1)/q^{m-t}\rfloor+1$ [15, Theorem 13], and $\mathcal{C}_{(q,m,q^{t}-2)}$ for $q>2$ [14, Theorem 15]. Recently, the dimension of $\mathcal{C}_{(q,m,q^{t}+b)}$ for $\lceil{m}/{2}\rceil\leq t<m$ and $0\leq b<q^{m-t}+\sum\limits_{i=1}^{\lfloor{t}/{s}\rfloor-1}q^{m+i(m-t)}$, and $\mathcal{C}_{(q,m,q^{t}-b)}$ for $\lceil{m}/{2}\rceil<t<m$ and $0\leq b<(q-1)\sum\limits_{i=1}^{m-t}q^{i}$ were provided in [21].
@@
In this paper, we determine the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/{3}\rfloor+1}.$ We also extend these results to BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ for $m\geq 4$ and positive integers $b$ and $\delta$ with $b+\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}+1$. In the existing literature, for $m\geq 4$, the dimension of $\mathcal{C}_{(q,m,\delta)}$ is known only for $\delta\in[2,q^{\lfloor(m+1)/{2}\rfloor+1}]$ and for some special cases. Moreover, it is clear that $q^{\lfloor(2m-1)/{3}\rfloor+1}\geq q^{\lfloor(m+1)/{2}\rfloor+1}\cdot q^{\lfloor(m-4)/{6}\rfloor}$. Therefore, our results provide the dimension of $\mathcal{C}_{(q,m,\delta)}$ for $\delta$ in a much larger range than previously known.
@@
Our results rely critically on the analysis of the distribution of coset leaders in $1,q^{\lfloor(2m-1)/3\rfloor+1}]$. Note that an integer $a\in[0,q^{m}-1]$ cannot be a coset leader if $q$ divides $a$. Therefore, to determine the number of coset leaders within a given range, we only need to count the integers that are neither divisible by $q$ nor coset leaders. We partitioned such integers in $\left[1,q^{\left\lfloor(2m-1)/3\right\rfloor+1}\right]$ into distinct classes with uniform representations by applying a one-to-one correspondence between the integer in $0,q^{m}-1]$ and length-$m$ sequences of non-negative integers less than $q$. This classification allows us to compute the number of integers in each class through a uniform method. Furthermore, we divide the interval $1,\delta-1]$ into subintervals based on the distribution of coset leaders, so that we can establish an equation to express the number of integers neither divisible by $q$ nor coset leaders in each subinterval. Consequently, we derive the exact count of coset leaders in $1,\delta-1]$ for each integer $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$. Combining this with our findings on the sizes of cosets, we ultimately determined the dimension of the code $\mathcal{C}_{(q,m,\delta)}$ for each integer $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$. On the other hand, the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ is indeed equal to the smallest coset leader not less than $\delta$. Therefore, it is not difficult to determine the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ based on our analysis of the distribution of coset leaders.
@@
$m$ is odd $2\leq\delta\leq q^{({m+1})/{2}}+1$ 1996 [48]
$m$ is even $2\leq\delta\leq 2q^{{m}/{2}}+1$ 1996 [48]
$m\geq 1$ $\delta=(q-l_{0})q^{m-l_{1}-1}-1$ with $0\leq l_{0}\leq q-2$ and $0\leq l_{1}\leq m-1$ 2015 [14]
$m$ is even $\delta=k(q^{{m}/{2}}+1)$ with $1\leq k\leq q-1$ 2015 [14]
@@
$m\geq 1$ $\delta=q^{t}+b$ with $1\leq b\leq\lfloor(q^{t}-1)/q^{m-t}\rfloor+1$ 2015 [15]
@@
$m=2$ $2\leq\delta\leq q^{m}-2$ 2017 [33]
$m>1$ $\delta=a({q^{m}-1})/{q-1}$ or $\delta=aq^{m-1}-1$ with $1\leq a\leq q-1$ 2020 [12]
@@
$m>1$ $\delta=q^{t}+b$ with $\lceil{m}/{2}\rceil\leq t<m$ and $0\leq b<q^{m-t}+\sum\limits_{i=1}^{\lfloor{t}/{s}\rfloor-1}q^{m+i(m-t)};$ $\delta=q^{t}-b$ with $\lceil{m}/{2}\rceil<t<m$ and $0\leq b<(q-1)\sum\limits_{i=1}^{m-t}q^{i}$ 2024 [21]
This paper is organized as follows. In Section II, we present some preliminaries that reveal the close relationship between $q$-cyclotomic cosets modulo $q^{m}-1$ and the dimension and Bose distance of BCH codes. Section III introduces some new results on $q$-cyclotomic cosets. These findings are then utilized to determine the dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor(2m-1)/{3}\rfloor+1}$ in Section IV. The result concerning the dimension of narrow-sense primitive BCH codes is subsequently extended to primitive BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ for some positive integers $\delta$ and $b$ with integers $\delta+b\leq q^{\lfloor(2m-1)/{3}\rfloor+1}+1$ in Section V. Moreover, we also apply the findings on $q$-cyclotomic cosets to obtain the Bose distance for primitive BCH codes in Section VI. We then provide examples of the BCH codes studied in this paper and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, in Section VII, as an illustration of our main results, we present the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $\delta=aq^{h+k}+b$ with integers $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, $a\in[1,q-1]$ and $b\in\left[1,q^{m-h-k}\right]$, where $h=\lfloor m/2\rfloor$. Finally, the conclusion of this paper is given in Section VIII.